Find the in slope intercept form of the passing through points with given coordiantes Find the equation of a line. (-5,-3) (-3,1)
Answer by stanbon(75887)

Find the in slope intercept form of the passing through points with given coordiantes Find the equation of a line. (-5,-3) (-3,1)
------------------
slope = (1--3)/(-3--5) = 4/2 = 2
---------------------------------------
Form: y = mx + b
1 = 2(-3) + b
b = 7
------------------
Equation:
y = 2x+7
======================
Cheers,
Stan H.